OpenStudy (anonymous):

Need reviewing help! Simplify: -3x^2y(x^2+2xy-5y^3)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

this is -3x^2y(x^2+2xy-5y^3) right?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

what is -3x^2y times x^2

OpenStudy (anonymous):

-6x^2y?

jimthompson5910 (jim_thompson5910):

the 2 in the x^2 does NOT multiply with the -3 to get -6

OpenStudy (anonymous):

Just kidding

OpenStudy (anonymous):

lol

jimthompson5910 (jim_thompson5910):

so what is it instead?

OpenStudy (anonymous):

wouldnt it still be the same?

jimthompson5910 (jim_thompson5910):

if you're stuck, then here's a similar example x^3 times x^5 = x^(3+5) = x^8

OpenStudy (anonymous):

or would it be -2x^2

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

-3x^2?

jimthompson5910 (jim_thompson5910):

x^2 times x^2 = ???

OpenStudy (anonymous):

x^2?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

x^2 itself means x squared anything square is the number multiplied by itself

jimthompson5910 (jim_thompson5910):

ex: 3^2 = 3*3 5^2 = 5*5

OpenStudy (anonymous):

so you add them instead of multiply?

jimthompson5910 (jim_thompson5910):

so... 5^2 times 5^2 = 5*5 times 5*5 = 5*5*5*5 = 5^4 shortcut: 5^2 times 5^2 = 5^(2+2) = 5^4

jimthompson5910 (jim_thompson5910):

replace the '5' with 'x' since you can do it with any base number, meaning that x^2 * x^2 = x^4

jimthompson5910 (jim_thompson5910):

so that brings us back to -3x^2y times x^2 = -3x^4y

jimthompson5910 (jim_thompson5910):

see how I got this?

OpenStudy (anonymous):

yeah, im sorry. i kept seeing it as 2y.

jimthompson5910 (jim_thompson5910):

oh no sry, -3x^2y is really -3*x^2*y

jimthompson5910 (jim_thompson5910):

i didn't feel like typing all those multiplication asterisks

OpenStudy (anonymous):

im super confused now. lol

jimthompson5910 (jim_thompson5910):

where at

jimthompson5910 (jim_thompson5910):

let's go back to the beginning of what the problem wants when you distribute this expression out -3x^2y(x^2+2xy-5y^3) you will get ( -3x^2y times x^2 ) + ( -3x^2y times 2xy ) + ( -3x^2y times -5y^3 ) basically you multiply the stuff on the outside, which is -3x^2y, by everything on the inside of the parenthesis and add

OpenStudy (anonymous):

Yeah, i did that its the rest im trying to do. Having a brain fart or whatever. :p

jimthompson5910 (jim_thompson5910):

ok tell me how far you got

jimthompson5910 (jim_thompson5910):

oh you got up to this ( -3x^2y times x^2 ) + ( -3x^2y times 2xy ) + ( -3x^2y times -5y^3 ) ?? if so, then you just need to multiply out each expression in the separate parenthesis groups

jimthompson5910 (jim_thompson5910):

do you know what I mean?

OpenStudy (anonymous):

Yeah

jimthompson5910 (jim_thompson5910):

ok tell me what you get then

OpenStudy (anonymous):

would the first one simp. be -2x^y

OpenStudy (anonymous):

ugh

jimthompson5910 (jim_thompson5910):

-3x^2y * x^2 -3(x^2*x^2)y -3x^(2+2)y -3x^4*y so, -3x^2y * x^2 turns into -3x^4*y

jimthompson5910 (jim_thompson5910):

notice how I grouped the terms with the same base

jimthompson5910 (jim_thompson5910):

then multiplied it out to simplify

OpenStudy (anonymous):

okay, that makes sense now. I just forgot how it worked.

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

glad it's clicking now though

jimthompson5910 (jim_thompson5910):

-3x^2y * 2xy turns into ????

OpenStudy (anonymous):

in this one would it be -6x^2y^2 or -6x^3y^2?

jimthompson5910 (jim_thompson5910):

the x^2 and the x multiply to x^3

OpenStudy (anonymous):

aalright.

jimthompson5910 (jim_thompson5910):

so it's the second one you got

OpenStudy (anonymous):

-6x^3y^2 and the third 15x^2y^4

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

good, you got the 3rd one correct

jimthompson5910 (jim_thompson5910):

so... ( -3x^2y times x^2 ) + ( -3x^2y times 2xy ) + ( -3x^2y times -5y^3 ) will turn into ( -3x^4y ) + (-6x^3y^2 ) + ( 15x^2y^4 ) and removing parenthesis gives -3x^4y - 6x^3y^2 + 15x^2y^4

OpenStudy (anonymous):

i got my answer as 6x^9y^5 is that right?

jimthompson5910 (jim_thompson5910):

you cannot combine any terms because there are no like terms

jimthompson5910 (jim_thompson5910):

besides, x^4 + x^3 is not x^7

OpenStudy (anonymous):

oh, so i just leave it after i take away the parenthesis?

jimthompson5910 (jim_thompson5910):

yeah you just leave something like x^4 + x^3 as it is...you cannot combine or simplify that at all

jimthompson5910 (jim_thompson5910):

if you had x^3 + x^3, that's a different story

OpenStudy (anonymous):

alrght(P:

jimthompson5910 (jim_thompson5910):

but you don't